# Recursive algorithms for computing Birkhoff interpolation polynomials

### v1

@@ 3 Main results
$p_{k-1}(x)$ $=-\left(\left[\begin{array}[]{c|cccc|c}0&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)&q_{k-1}(x)\\ \hline\cr y_{\beta_{1},\alpha_{1}}&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ y_{\beta_{2},\alpha_{2}}&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots&\vdots\\ y_{\beta_{k-1},\alpha_{k-1}}&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \hline\cr y_{\beta_{k},\alpha_{k}}&q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/V_{k-1}\right)$
$\triangleq-\left(\left[\begin{array}[]{ccc}0&Q_{k-2}(x)&q_{k-1}(x)\\ Y_{k-1}&V_{k-2}&B\\ y_{\beta_{k},\alpha_{k}}&Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/\left(\begin{array}[]{cc}V_{k-2}&B\\ Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right)\right).$
@@
$p_{k-1}(x)$ $=-\left(\left(\begin{array}[]{cc}0&Q_{k-2}(x)\\ Y_{k-1}&V_{k-2}\\ \end{array}\right)/V_{k-2}\right)$
$~~~~~+\left(\left(\begin{array}[]{cc}Q_{k-2}(x)&q_{k-1}(x)\\ V_{k-2}&B\\ \end{array}\right)/V_{k-2}\right)\left(\left(\begin{array}[]{cc}V_{k-2}&B\\ Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right)/V_{k-2}\right)^{-1}\left(\left(\begin{array}[]{cc}Y_{k-1}&V_{k-2}\\ y_{\beta_{k},\alpha_{k}}&Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})\\ \end{array}\right)/V_{k-2}\right).$
@@
$g_{j,k-1}(x)=g_{j-1,k-1}(x)-\frac{g_{j-1,k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})}{g_{j-1,j}^{(\alpha_{j+1})}(x_{\beta_{j+1}})}g_{j-1,j}(x),~j=0,1,\dots,k-2.$ (34)
@@
$g_{j,k-1}(x)$ $=\left(\left[\begin{array}[]{c|cccc|c}q_{k-1}(x)&q_{0}(x)&q_{1}(x)&\dots&q_{j-1}(x)&q_{j}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{j-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{j}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{j-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{j}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots&\vdots\\ q_{k-1}^{(\alpha_{j})}(x_{\beta_{j}})&q_{0}^{(\alpha_{j})}(x_{\beta_{j}})&q_{1}^{(\alpha_{j})}(x_{\beta_{j}})&\dots&q_{j-1}^{(\alpha_{j})}(x_{\beta_{j}})&q_{j}^{(\alpha_{j})}(x_{\beta_{j}})\\ \hline\cr q_{k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{0}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&\dots&q_{j-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j+1})}(x_{\beta_{j+1}})\\ \end{array}\right]/V_{j}\right)$
$\triangleq-\left(\left[\begin{array}[]{ccc}q_{k-1}(x)&Q_{j-1}(x)&q_{j}(x)\\ \tilde{B_{1}}&V_{j-1}&\tilde{B}\\ q_{k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&Q_{j-1}^{\alpha_{j+1}}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j}+1)}(x_{\beta_{j+1}})\\ \end{array}\right]/\left(\begin{array}[]{cc}V_{j-1}&\tilde{B}\\ Q_{j-1}^{\alpha_{j+1}}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j}+1)}(x_{\beta_{j+1}})\\ \end{array}\right)\right).$
@@ 4.1 Algorithm 1
Now we present an algorithm to compute the monomial basis, Newton-type basis, and interpolation polynomial to the interpolation conditions specified in (2). In the algorithm below, $\{g_{-1,i}(x),i=0,1,\dots,N-1\}$ are recursively updated.
@@
Since $\textup{span}\{g_{-1,0}(x),g_{-1,1}(x),\dots,g_{-1,k-1}(x)\}=\textup{span}\{g_{-1,0}(x),g_{0,1}(x),\dots,g_{k-2,k-1}(x)\}$, it follows that $\{g_{-1,0}(x),g_{-1,1}(x),\dots,g_{-1,k-1}(x)\}$ constitutes a set of strongly proper monomial basis. ∎
Example 1 We consider the same example as in [17]. Let $x_{0}=1,x_{1}=2,x_{2}=3$, $y_{0,0}=5,y_{1,1}=6,y_{1,2}=4,y_{2,2}=7$. By the interpolation conditions, we can get the incidence matrix
$E=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&1&0\\ 0&0&1&0\\ \end{array}\right),$
and this incidence matrix satisfies the Polya’s conditions. In Algorithm 1, the interpolation conditions are sorted into NN-DOS as follows
@@
According to Algorithm 1, we can obtain
Table 1: The auxiliary polynomials
@@
In Table 1,
$g_{-1,0}(x_{0})=1,\ g_{0,1}^{(1)}(x_{1})=1,\ g_{1,2}^{(2)}(x_{1})=2,\ g_{2,3}^{(2)}(x_{2})=6.$
and $\{1,x-1,x^{2}-4x+3,x^{3}-6x^{2}+12x-7\}$ is a Newton-type basis and the interpolation polynomial is
@@
Example 2 Let $x_{0}=-1,x_{1}=0,x_{2}=1$, $y_{0,0}=2,y_{1,1}=4,y_{2,0}=6,y_{2,1}=8$. Based on the above conditions, we can derive the incidence matrix
$E=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 1&1&0&0\\ \end{array}\right),$
and this incidence matrix does satisfy the Polya’s conditions. By Definition 2,
@@
According to Algorithm 1, we can derive
Table 2: The auxiliary polynomials
@@
in which $\{g_{-1,0}(x),g_{-1,1}(x),g_{-1,2}(x),g_{-1,3}(x)\}=\{1,x,x^{3},x^{4}\}$ is a strongly proper monomial basis. It is easy to verify that
@@
In fact, the Vandermonde matrix corresponding to the interpolation problem is
[…]
That is, the set $\{1,x+1,x^{3}-x,x^{4}-1\}$ constitutes the Newton-type basis with respect to the given interpolation conditions. The interpolation polynomial computed by the recursive algorithm is
@@ 4.2 Algorithm 2
By computing the Vandermonde determinant, we can verify that the set $\{1,x,x^{2},x^{3}\}$ also constitutes a strongly proper monomial basis for Example 2. This implies that the interpolation polynomial produced by Algorithm 1 is not necessarily of minimal degree. Furthermore, Algorithm 1 is restricted to cases where the interpolation conditions consist of monomial differential operators and evaluation functionals. We enhance Algorithm 1, making it capable of handling interpolation conditions formed by the composition of differential polynomials and evaluation functionals. Moreover, Algorithm 2 offers the additional benefit of degree reduction in specific examples by reordering the interpolation conditions.
In Algorithm 2, we let $L_{1},L_{2},\dots,L_{N}$ denote the interpolation conditions, where every $L_{i},i=1,2,\dots,N$ is the composite operation of an evaluation functional and a differential polynomial, and we sort them in ascending order of the highest derivative order they involve. Namely, if we let $\alpha_{i}$ be the highest-order derivative of $L_{i}$, then $\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{N}$. For example, suppose the interpolation conditions at $x_{0}$ are $\delta_{x_{0}}\circ(D_{x}),\delta_{x_{0}}\circ(1+D_{x}^{2})$, and the interpolation conditions at $x_{1}$ are $\delta_{x_{1}}\circ(1+D_{x}),\delta_{x_{1}}\circ(D_{x}^{2}+D_{x}^{3})$, we then sort the interpolation conditions into the sequence:
$L_{1}=\delta_{x_{0}}\circ(D_{x}),\ L_{2}=\delta_{x_{1}}\circ(1+D_{x}),\ L_{3}=\delta_{x_{0}}\circ(1+D_{x}^{2}),\ L_{4}=\delta_{x_{1}}\circ(D_{x}^{2}+D_{x}^{3}).$
@@
Table 3: The auxiliary polynomials
@@ 5 Conclusions
Due to the discontinuity of the derivative conditions at the interpolation nodes in the Birkhoff interpolation problem, the set of all polynomials satisfying the homogeneous interpolation conditions fails to form an ideal. Consequently, the computation of the corresponding interpolation basis and the interpolating polynomial becomes more complex. This paper generalizes the GRPIA and presents two recursive algorithms for computing the strongly proper interpolation basis, the Newton-type basis, and the interpolating polynomial of the Birkhoff interpolation problems. For the case where the interpolation conditions consist of monomial differential operators and evaluation functionals, we first reorder the interpolation conditions to obtain an NN-DOS. Then, using the theory of Schur complement, we derive formulas for the corresponding Newton basis and the interpolating polynomial recursively. In contrast to existing algorithms that compute the Birkhoff interpolation basis using Gaussian elimination, our recursive approach achieves a substantial reduction in computational cost. Simultaneously, Algorithm 1 incorporates a core judgment condition (Line 18 of Algorithm 1) to ensure the non-singularity of the Vandermonde matrix for the Birkhoff interpolation problem. It is possible that in some numerical examples, Algorithm 1 may not yield a minimal-degree interpolation basis. To obtain lower-degree interpolation basis, we develop Algorithm 2 as an improved version. Furthermore, Algorithm 2 is capable of solving more general Birkhoff interpolation problems, where the conditions are composed of differential polynomials and evaluation functionals.
[…]
Author Contributions All authors contributed equally to the paper.
[…]
Funding Funding was provided by the Department of Education of Jilin Province (JJKH20250468KJ), China.
Declarations
[…]
Conflicts of interest The authors declare that they have no conflict of interest.
[…]
Competing interests The authors declare no competing interests.

### v2

@@ 3 Main results
$\begin{aligned} p_{k-1}(x)&=-\left(\left[\begin{array}[]{c|cccc|c}0&q_{0}(x)&q_{1}(x)&\dots&q_{k-2}(x)&q_{k-1}(x)\\ \hline\cr y_{\beta_{1},\alpha_{1}}&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{k-2}^{(\alpha_{1})}(x_{\beta_{1}})&q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})\\ y_{\beta_{2},\alpha_{2}}&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{k-2}^{(\alpha_{2})}(x_{\beta_{2}})&q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots&\vdots\\ y_{\beta_{k-1},\alpha_{k-1}}&q_{0}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&\dots&q_{k-2}^{(\alpha_{k-1})}(x_{\beta_{k-1}})&q_{k-1}^{(\alpha_{k-1})}(x_{\beta_{k-1}})\\ \hline\cr y_{\beta_{k},\alpha_{k}}&q_{0}^{(\alpha_{k})}(x_{\beta_{k}})&q_{1}^{(\alpha_{k})}(x_{\beta_{k}})&\dots&q_{k-2}^{(\alpha_{k})}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/V_{k-1}\right)\\ &\triangleq-\left(\left[\begin{array}[]{ccc}0&Q_{k-2}(x)&q_{k-1}(x)\\ Y_{k-1}&V_{k-2}&B\\ y_{\beta_{k},\alpha_{k}}&Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right]/\left(\begin{array}[]{cc}V_{k-2}&B\\ Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right)\right).\end{aligned}$
@@
$p_{k-1}(x)$ $=-\left(\left(\begin{array}[]{cc}0&Q_{k-2}(x)\\ Y_{k-1}&V_{k-2}\\ \end{array}\right)/V_{k-2}\right)+\left(\left(\begin{array}[]{cc}Q_{k-2}(x)&q_{k-1}(x)\\ V_{k-2}&B\\ \end{array}\right)/V_{k-2}\right)$
$~~~~~~~~\left(\left(\begin{array}[]{cc}V_{k-2}&B\\ Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})&q_{k-1}^{(\alpha_{k})}(x_{\beta_{k}})\\ \end{array}\right)/V_{k-2}\right)^{-1}\left(\left(\begin{array}[]{cc}Y_{k-1}&V_{k-2}\\ y_{\beta_{k},\alpha_{k}}&Q_{k-2}^{\alpha_{k}}(x_{\beta_{k}})\\ \end{array}\right)/V_{k-2}\right).$
@@
$g_{j,k-1}(x)=g_{j-1,k-1}(x)-\frac{g_{j-1,k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})}{g_{j-1,j}^{(\alpha_{j+1})}(x_{\beta_{j+1}})}g_{j-1,j}(x),~j=0,1,\dots,k-2.$
@@
$\begin{aligned} g_{j,k-1}(x)&=\left(\left[\begin{array}[]{c|cccc|c}q_{k-1}(x)&q_{0}(x)&q_{1}(x)&\dots&q_{j-1}(x)&q_{j}(x)\\ \hline\cr q_{k-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{0}^{(\alpha_{1})}(x_{\beta_{1}})&q_{1}^{(\alpha_{1})}(x_{\beta_{1}})&\dots&q_{j-1}^{(\alpha_{1})}(x_{\beta_{1}})&q_{j}^{(\alpha_{1})}(x_{\beta_{1}})\\ q_{k-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{0}^{(\alpha_{2})}(x_{\beta_{2}})&q_{1}^{(\alpha_{2})}(x_{\beta_{2}})&\dots&q_{j-1}^{(\alpha_{2})}(x_{\beta_{2}})&q_{j}^{(\alpha_{2})}(x_{\beta_{2}})\\ \vdots&\vdots&\vdots&&\vdots&\vdots\\ q_{k-1}^{(\alpha_{j})}(x_{\beta_{j}})&q_{0}^{(\alpha_{j})}(x_{\beta_{j}})&q_{1}^{(\alpha_{j})}(x_{\beta_{j}})&\dots&q_{j-1}^{(\alpha_{j})}(x_{\beta_{j}})&q_{j}^{(\alpha_{j})}(x_{\beta_{j}})\\ \hline\cr q_{k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{0}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&\dots&q_{j-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j+1})}(x_{\beta_{j+1}})\\ \end{array}\right]/V_{j}\right)\\ &\triangleq-\left(\left[\begin{array}[]{ccc}q_{k-1}(x)&Q_{j-1}(x)&q_{j}(x)\\ \tilde{B_{1}}&V_{j-1}&\tilde{B}\\ q_{k-1}^{(\alpha_{j+1})}(x_{\beta_{j+1}})&Q_{j-1}^{\alpha_{j+1}}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j}+1)}(x_{\beta_{j+1}})\\ \end{array}\right]/\left(\begin{array}[]{cc}V_{j-1}&\tilde{B}\\ Q_{j-1}^{\alpha_{j+1}}(x_{\beta_{j+1}})&q_{j}^{(\alpha_{j}+1)}(x_{\beta_{j+1}})\\ \end{array}\right)\right).\end{aligned}$
@@ 4.1 Algorithm 1
[…]
@@
Since
$\textup{span}\{g_{-1,0}(x),g_{-1,1}(x),\dots,g_{-1,k-1}(x)\}=\textup{span}\{g_{-1,0}(x),g_{0,1}(x),\dots,g_{k-2,k-1}(x)\},$
it follows that $\{g_{-1,0}(x),g_{-1,1}(x),\dots,g_{-1,k-1}(x)\}$ constitutes a set of strongly proper monomial basis. ∎
[…]
Theorem 10.
For any fixed $k=2,3,\dots,N$, suppose that the computation of the first k2k-2 columns has already been completed. Then, for $j=0,1,\dots,k-2$, the combination coefficients of $g_{j,k-1}(x)$ are precisely the elimination coefficients obtained when performing column elimination on the Vandermonde matrix constructed from the basis
$\{g_{-1,0}(x),g_{0,1}(x),\dots,g_{k-3,k-2}(x),g_{j-1,k-1}(x)\}$
under the first $k$ interpolation conditions—specifically, elimination is carried out using the nonzero diagonal element in the $j$-th column to annihilate $g_{j-1,k-1}^{(\alpha_{j+1})}(x_{\beta{j+1}})$ to zero.
Example 1 We consider the same example as in [16]. Let $x_{0}=1,x_{1}=2,x_{2}=3$, $y_{0,0}=5,y_{1,1}=6,y_{1,2}=4,y_{2,2}=7$. The interpolation conditions are sorted into NN-DOS as follows
@@
According to Algorithm 1, all auxiliary polynomials are listed in Table 1.
Table 1: The auxiliary polynomials.
@@
We can verify that the Vandermonde matrix takes the following form
[…]
The elements in the strictly lower triangular part of the matrix need not be calculated. $\{1,x-1,x^{2}-4x+3,x^{3}-6x^{2}+12x-7\}$ is a Newton-type basis and the interpolation polynomial is
@@
Remark 11.
Theorem 10 indicates that the construction of the auxiliary polynomials in Algorithm 1 implicitly embodies the idea of Gaussian elimination. Compared with existing works that apply Gaussian elimination to solve Birkhoff interpolation problems [7], the present approach differs mainly in the following two aspects: 1) During the iterative procedure of Algorithm 1, when computing the $k$-th Newton interpolation basis, $k=1,2,\dots,N$, only the first $k$ interpolation conditions are evaluated on the auxiliary polynomial and processed by Gaussian elimination, without considering all $N$ conditions. As illustrated in Example 1, the entries below the main diagonal of the resulting Vandermonde matrix are never accessed during the algorithm, which reduces computational cost. 2) In the process of sequentially computing all auxiliary polynomials in the (k1)(k-1)-th column, once a later polynomial is computed using the previous one, the earlier polynomials need not be retained. Ultimately, only the diagonal term $g_{k-2,k-1}(x)$ needs to be preserved. This significantly reduces storage space.
Remark 12.
While both the proposed algorithm and the method of [16] are generalizations of GRPIA [12] and employ the Schur complement theory within a univariate Birkhoff interpolation framework, they are rooted in fundamentally different principles. In [16], the authors address the Birkhoff interpolation problem by constructing high-degree interpolation bases that admit factorization, ultimately yielding a simplified version of the RHBPIA algorithm. The proposed Algorithm 1 introduces a criterion to ensure the well-posedness of the Birkhoff interpolation problem, thereby yielding lower-degree Newton bases and interpolation polynomials. We further propose Algorithm 2 to address a more general Birkhoff interpolation problem whose conditions are composed of evaluation functionals and differential polynomials, extending beyond those involving only partial derivatives. This extension represents another key difference from the RHBPIA method.
Example 2 Let $x_{0}=-1,x_{1}=0,x_{2}=1$, $y_{0,0}=2,y_{1,1}=4,y_{2,0}=6,y_{2,1}=8$. By Definition 1,
@@
According to Algorithm 1, the auxiliary polynomials are shown in the following table.
Table 2: The auxiliary polynomials.
@@
$\{g_{-1,0}(x),g_{-1,1}(x),g_{-1,2}(x),g_{-1,3}(x)\}=\{1,x,x^{3},x^{4}\}$ is a strongly proper monomial basis. It is easy to verify that
@@
The set $\{1,x+1,x^{3}-x,x^{4}-1\}$ constitutes the Newton-type basis with respect to the given interpolation conditions. The interpolation polynomial computed by the recursive algorithm is
@@ 4.2 Algorithm 2
By computing the Vandermonde determinant, we can verify that the set $\{1,x,x^{2},x^{3}\}$ also constitutes a proper monomial basis for Example 2. This implies that the interpolation polynomial produced by Algorithm 1 is not necessarily of minimal degree. Furthermore, Algorithm 1 is restricted to cases where the interpolation conditions consist of monomial differential operators and evaluation functionals. We enhance Algorithm 1, making it capable of handling interpolation conditions formed by the composition of differential polynomials and evaluation functionals. Moreover, Algorithm 2 offers the additional benefit of degree reduction in specific examples by reordering the interpolation conditions.
In Algorithm 2, we let $L_{1},L_{2},\dots,L_{N}$ denote the interpolation conditions, where every $L_{i},i=1,2,\dots,N$ is the composite operation of an evaluation functional and a differential polynomial, and we sort them in ascending order of the highest derivative order they involve. Namely, if we let $\alpha_{i}$ be the highest-order derivative of $L_{i}$, then $\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{N}$. For example, suppose the interpolation conditions at $x_{0}$ are $\delta_{x_{0}}\circ D_{x},\delta_{x_{0}}\circ(1+D_{x}^{2})$, and the interpolation conditions at $x_{1}$ are $\delta_{x_{1}}\circ(1+D_{x}),\delta_{x_{1}}\circ(D_{x}^{2}+D_{x}^{3})$, we then sort the interpolation conditions into the sequence:
@@
Table 3: The auxiliary polynomials.
@@ 5 Conclusions
Due to the discontinuity of the derivative conditions at the interpolation nodes in the Birkhoff interpolation problem, the set of all polynomials satisfying the homogeneous interpolation conditions fails to form an ideal. Consequently, the computation of the corresponding interpolation basis and the interpolating polynomial becomes more complex. This paper generalizes the GRPIA and presents two recursive algorithms for computing the strongly proper interpolation basis, the Newton-type basis, and the interpolating polynomial of the Birkhoff interpolation problems. For the case where the interpolation conditions consist of monomial differential operators and evaluation functionals, we first reorder the interpolation conditions to obtain an NN-DOS. Then, using the theory of Schur complement, we derive formulas for the corresponding Newton basis and the interpolating polynomial recursively. In contrast to existing algorithms that compute the Birkhoff interpolation basis using Gaussian elimination, our recursive approach achieves a substantial reduction in computational cost and storage space. Simultaneously, Algorithm 1 incorporates a core judgment condition (Line 18 of Algorithm 1) to ensure the non-singularity of the Vandermonde matrix for the Birkhoff interpolation problem. It is possible that in some numerical examples, Algorithm 1 may not yield a minimal-degree interpolation basis. To obtain lower-degree interpolation basis, we develop Algorithm 2 as an improved version. Furthermore, Algorithm 2 is capable of solving more general Birkhoff interpolation problems, where the conditions are composed of differential polynomials and evaluation functionals.
Acknowledgments
This research was supported by the Department of Education of Jilin Province (JJKH20250468KJ), China.